OpenStudy (abbycross167):

can someone please help me answer this question? I could really use some help... Question~ Solve the systems using the elimination method. y = 2x x + y = 3

OpenStudy (wolf1728):

Abby I'd say the first thing to do would be to put the first equation into proper form.

OpenStudy (abbycross167):

what do you mean?

OpenStudy (wolf1728):

Okay I'll help you with that. Basically when you are solving for 2 unknowns you want both equations set up similarly. So, we'll take y = 2x and change it to -2x + y -0

OpenStudy (wolf1728):

Sorry -2x + y = 0

OpenStudy (anonymous):

X=1 and y=2 I plugged in random numbers and see if they worked and these did...:))

OpenStudy (anonymous):

Oh just kidding no they don't hold on

OpenStudy (wolf1728):

Sydney - just picking and choosing MIGHT work sometimes but not always

OpenStudy (anonymous):

I always do... And I have an a... I've never actually missed one of these questions, sometimes it takes forever tho

OpenStudy (abbycross167):

@wolf1728 could you show me the steps to answer this question? please?

OpenStudy (wolf1728):

Yes, but I expect you to type some things every now and then

OpenStudy (wolf1728):

Okay we now have the two equations -2x + y = 0 x + y = 3

OpenStudy (abbycross167):

yes sir/ma'am ofc! Thank you very much!

OpenStudy (abbycross167):

Multiply the 2nd equation by 3(or would it be 2?) so that the y’s have opposite coefficients?

OpenStudy (wolf1728):

I'm sorry We should multiply the first equation by -1 then we have 2x - y = 0 x + y = 3 Now it's time to add the 2 equations.

OpenStudy (wolf1728):

Adding those we have 2x plus x = ?

OpenStudy (wolf1728):

abby 2x + x = ?

OpenStudy (abbycross167):

0? I'm sorry it takes me a minute to figure it out...

OpenStudy (wolf1728):

abby I'm just asking what is 2x + x

OpenStudy (abbycross167):

what do you mean? do you want to know what it equals?

OpenStudy (wolf1728):

Yes abby what is 2x + x?

OpenStudy (abbycross167):

0?

OpenStudy (wolf1728):

No If you have two x's and 1 x how many x'sdo you have?

OpenStudy (abbycross167):

2?

OpenStudy (wolf1728):

Basically what is 2+1?

OpenStudy (abbycross167):

3

OpenStudy (wolf1728):

Okay So we are in the process of adding the 2 equations 2x - y = 0 x + y = 3 2x + x = 3x -y + y = 0 and 3 +0 =3 So, we now have 3x = 3 What does x equal?

OpenStudy (wolf1728):

What times 3 equals 3?

OpenStudy (abbycross167):

1

OpenStudy (wolf1728):

Yes so x =1 To solve for y, let's use equation 2 x + y = 3 We know x = 1 So, 1 + y = 3 What is y?

OpenStudy (abbycross167):

2?

OpenStudy (wolf1728):

Yes, so now you know x =1 and y = 2

OpenStudy (abbycross167):

oh ok and that's the answer to the question?

OpenStudy (wolf1728):

Yes, that is the answer

OpenStudy (abbycross167):

ok great! thank you soo much @wolf1728 you're awesome! I'm going to write out the steps... would you mind looking at them after I wright them out to make sure I did it right?

OpenStudy (wolf1728):

I must get going now

OpenStudy (abbycross167):

ok well thank you!!